Project

General

Profile

NDOS Commissioning Meeting 2011/10/31

Attendance: Jon, Athans, Minerba, Raphael, Leon, Mat, Enrique, Martin, Kurt

Agenda:

  1. Beam & Detector Status
  2. DAQ/Electronics Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & Detector Update

  • Mat:
    • cooled channels remain unchanged; reshuffled APDs seem to be generally ok, but a few are being masked off because they are slightly above the cut. Will look in more detail.
    • leaky parylene APD is still in the readout and was thru the weekend. But Leon notes that it doesn't seem to be producing real hits.
    • new mounting hardware will be arriving at FNAL later this weekend, and will attempt dry assembly (no APD). Should not require detector downtime. Leon: will new parts fit in old box?

DAQ/Electronics Status & Plans

  • from Peter: "... we're most of the way through the updates that came up last week, except for Kurt's ConfigManager update that will have to drag along a few other packages. I would like to get those in today, and then setup a frozen release today or tomorrow."

DCS Status & Plans

  • Martin:
    • updated DCS code to work with additional DCMs (3-2, 3-3) and it all seems to be working nicely.
    • lowered the set point on temperature of FEB 2-3-11 and the drive current seems to be behaving better.

AOB

  • would like to move to only one meeting per week; Jon suggested Monday mornings, Leon prefers Thursday afternoon. Thursday it is!